OpenStudy (anonymous):

Rationalize the denominator and simplify 2/(4- sqrt5)

OpenStudy (anonymous):

multiply both numerator and denominator with 4+sqrt5

OpenStudy (ash2326):

@jmallis in previous question. I had solved by rationalizing the denominator . Can you try it here also. I'll give you a hint use \(4+\sqrt{5}\) to rationalize the denominator

OpenStudy (anonymous):

\[\frac{2}{4-\sqrt{5}}\times \frac{4+\sqrt{5}}{2+\sqrt{5}}\]is a start. multiply top and bottom by conjugate of the bottom

OpenStudy (anonymous):

8+2sqrt5/11 ?

OpenStudy (ash2326):

Yeah you're right. Great work:D